# Make Me Happier: Evoking Emotions Through Image Diffusion Models

### v3

@@ 6 Discussion
(c) Image pairs used in the entire experiment are drawn from the 2,016 result pairs generated by our model and five other SOTA methods. Participants view randomly sampled pairs from these 2,016 pairs, and all trials are presented in a random order. The order of the two images for selection is also randomized. An example trial in the real experiment is illustrated in Figure S6(d).
Figure S6: MTurk Experiment. (a) Schematic of MTurk experiment. Participants are presented with a set of images and a target emotion. They must choose between two images, one generated by our model and the other by five state-of-the-art methods, selecting the one that more effectively evokes the target emotion. (b) Instructions of the MTurk experiment. (c) Dummy trial example. We chose six image pairs with prominent emotional distinctions as benchmarks to assess participants’ comprehension of the task and their attentiveness to the experiment. (d) Real trial example.
@@
We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)Turk (2012) (Online). We recruited 417 participants, each undergoing 100 trials. Additionally, we conducted in-lab experiments, involving 10 volunteers. Similarly, each participant reviewed 100 trials. The consistency between online and in-lab experimental results demonstrates the superiority of our approach in producing results that evoke the target emotion compared to other state-of-the-art methods.
Figure S7: Preference of Our Method Over SOTA Methods. We conducted both online experiments (417 participants) and in-lab experiments (10 participants). The left column represents the online results, while the right column displays the in-lab results. The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions. Chance is 50% in the red dotted line. Error bars are standard errors.
Figure S7 illustrates the preference of our method over other state-of-the-art methods in both online and in-lab experiments. From the overall results (Row 1), the in-lab experiment demonstrates a higher preference compared to the online experiment. This suggests that our method’s effectiveness in evoking the target emotion is more pronounced in a laboratory setting. Note that our in-lab and online experiments had identical setups. The observed difference may stem from the controlled environment provided by the lab conditions, where participants were more focused, reducing randomness and facilitating a more accurate assessment of emotion generation. This is further supported by the dummy trial success rates, with a 100% pass rate in the in-lab experiment compared to only 27% in the online experiment. Again, we emphasize that we only use the online data for result analysis, where the participants passed the dummy tests with less than 1/6 error rates.
Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-TransferGatys et al. (2015) (NST) and CLIP-StylerKwon, Ye (2022) (Csty) exhibit an advantage in generating negative results, but struggle with positive outcomes. Additionally, our EmoEditor surpasses Color-transferPitié et al. (2007) (CT) and Ip2pBrooks et al. (2023) in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions. This indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones. Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable; however, our approach allows for end-to-end training and testing, avoiding the complexity of chaining large models.
@@
In addition, we show the confusion matrix of different methods for both online and in-lab experiments in Figure S8 and S9. The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories. From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer Gatys et al. (2015) and CLIP-Styler Kwon, Ye (2022) more in the manipulation of emotions from positive to negative. This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image. In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images compared to these two approaches.
Figure S8: Confusion matrix of different methods (Online). The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs. The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods. The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods. Average preferences for our EmoEditor and other methods are indicated above each matrix. The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion. Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
@@
Figure S9: Confusion matrix of different methods (In-lab). The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs. The top-left matrix presents the average overall results of our EmoEditor compared to the other five SOTA methods. The remaining five matrices show individual comparisons between our EmoEditor and each of the five SOTA methods. Average preferences for our EmoEditor and other methods are indicated above each matrix. The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion. Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
@@
Figure S10 illustrates the visualizations of images generated by all methods across various scenarios. Color-Transfer Pitié et al. (2007) (CT) proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image. For instance, in Row 1, Column 2, it enhances brightness but fails to evoke amusement as it does not alter elements causing sadness, such as the crying woman. Neural-Style-Transfer Gatys et al. (2015) (NST) heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image. For example, in Row 4, Column 3, it generates multiple differently colored blocks but fails to evoke the emotion of sadness. CLIPstyler Kwon, Ye (2022) (Csty) introduces inexplicable textures. Ip2p Brooks et al. (2023), due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities, as seen in Column 5.
Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion, as evidenced by Row 1, Column 6, where it not only fails to change the crying woman but amplifies the crying details. In addition, it often greatly changes the scene of the source image to achieve the goal of changing the emotion, such as Row 2, Column 6, it produces a building that looks very technological but quite different from the source image.
In contrast, our EmoEditor, requiring only the target emotion and source image inputs, can generate highly creative images that evoke the target emotion while striving to maintain scene structure and semantic coherence. For example, in Row 1, Column 7, it changes the woman’s pose while retaining the main scene and character from the source image, generating a cartoon-like image that aligns with amusement. In Row 2, Column 7, our EmoEditor generates technological results that make people feel awe while almost maintaining most of the structure of the source image.
Figure S10: Visualization of Generated Images from Different Methods. The target emotion is highlighted in red and the source image is framed with green.
S3.3 Ablation Study of Different Iterations
Due to the abstract and complex nature of emotions, we propose an iterative emotion inference approach. The maximum number of iterations significantly impacts the quality of images generated by the model. To assess this effect, we vary the maximum number of iterations from 1 to 30 and evaluate the results using four metrics.
Figure S11 presents our experimental findings. The Emotion Matching Ratio (EMR) shows a steady increase from 0% to approximately 50% over 30 iterations, indicating continuous improvement in the model’s ability to generate images that evoke the target emotions.
The Emotion Similarity Ratio (ESR) rises rapidly in the initial iterations, stabilizing around 90%. This suggests that the model quickly becomes proficient at producing images that closely match the target emotional content, with minimal improvements after the initial iterations.
The Emotion-Neutral Region Deviation (ENRD) also increases rapidly at first, plateauing around 22.5. This pattern indicates that the deviation in neutral regions of the generated images becomes consistent after a few iterations, reflecting the model’s stability in preserving non-emotional aspects.
The Edge Structure Similarity (ESS) follows a similar trend, rising quickly to around 16 and then stabilizing. This stability in ESS reflects the model’s ability to maintain structural coherence and semantic consistency between the source and generated images after the initial iterations.
In summary, these metrics demonstrate that the iterative process enhances the model’s performance significantly in the initial stages, reaching a stable state of high quality and consistency in generating emotionally evocative images.
Figure S11: Performance Metrics Across Different Iterations.
@@
Figure S12 shows some failure cases. Our EmoEditor encounters difficulties in handling fine details of individuals, especially when their faces in the images are small. For example, in Row 1, although our EmoEditor attempts to make the boy appear sad, the boy’s face seems somewhat twisted and distorted.
Moreover, significant disparities between source images and target emotions present obstacles, making it difficult to generate results aligning with the target emotion while preserving the structure of the source image. For example, in Row 2, the source image features a tiger, evoking the emotion of anger. In our EmoEditor’s attempt to transform it into awe, although the image’s structure is retained, it is altered to depict steep coastal rocks, changing the semantic content of the image.
Figure S12: Failure Cases. The target emotion is highlighted in red and the source image is framed with green.

### v4

@@ 6 Discussion
(c) Image pairs used in the entire experiment are drawn from the 2,016 result pairs generated by our model and other SOTA methods.
Participants view randomly sampled pairs from these 2,016 pairs, and all trials are presented in a random order. The order of the two images for selection is also randomized. An example trial in the real experiment is illustrated in Figure S6(d).
Figure S6: MTurk Experiment. (a) Schematic of MTurk experiment. Participants are presented with a set of images and a target emotion. They must choose between two images, one generated by our model and the other by 9 state-of-the-art methods, selecting the one that more effectively evokes the target emotion. (b) Instructions of the MTurk experiment. (c) Dummy trial example. We chose six image pairs with prominent emotional distinctions as benchmarks to assess participants’ comprehension of the task and their attentiveness to the experiment. (d) Real trial example.
S2.4 Evaluation Metrics: Structural Score
Emotion-evoked image generation needs to be evaluated from both emotional induction and structural preservation perspectives. Traditional metrics, such as SSIM, are inadequate for structural preservation evaluation because they do not consider emotional cues. The regions in the image that evoke emotions are called Emotion-Evoking Regions ($R_{emo}$). These areas need to be altered to change the source emotion and induce the target emotion. Conversely, some regions in the image lack significant emotional cues and are referred to as Emotion-Neutral Regions ($R_{neu}$). These regions should be preserved as much as possible during the editing process.
Grad-CAM [24] can visualize which parts of the image most influence the model’s decision. Therefore, we use Grad-CAM with the emotion predictor $\mathcal{P}$ to identify Emotion-Evoking Regions ($R_{emo}$) in the source images. Figure S7 shows the visualization of the Emotion-Evoking and Emotion-Neutral Regions. Using Grad-CAM [24] with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify Emotion-Evoking Regions (valued at 1) and Emotion-Neutral Regions (valued at 0) on source images.
To evaluate changes, we calculate pixel-level L1 differences between the source and generated images for these regions, denoted as $L_{emo}$ and $L_{neu}$. The structural score is: $S_{str}=L_{emo}/(L_{emo}+L_{neu}+\epsilon)$ where $\epsilon$ is a small constant to avoid division by zero. This score reflects the proportion of changes in $R_{emo}$ relative to all changed pixels, with higher values indicating better preservation of neutral areas.
Figure S7: Visualization Explanation of the Emotion-Evoking and Emotion-Neutral Regions. The source image is framed in green. Using Grad-CAM [24] with $\mathcal{P}$, we binarize Grad-CAM maps at a 0.5 threshold to identify Emotion-Evoking Regions (valued at 1) and Emotion-Neutral Regions (valued at 0) on source images.
Considering that Grad-CAM may be inaccurate or unstable due to the limitations of the emotion predictor $\mathcal{P}$, we invited 46 participants to annotate the emotion-evoking regions of 504 source images. Each worker annotated 50 trials, resulting in a total of 2,300 trials. Additionally, we included 6 dummy trials, which were the same as those used in our human psychophysics experiment (Figure S6(c)), to ensure annotation quality. Results from workers who failed more than once on the dummy trials were discarded. Ultimately, each image was annotated by at least three individuals. We then averaged all the annotations and binarized the results to generate a human-annotated weight map. Figure S8 shows example annotations and a comparison with Grad-CAM.
Figure S8: Emotion-Evoking Region Annotated by Human. Each image has been annotated by at least three human annotators who identified the emotion-evoking regions. We averaged all the annotations and binarized the result to generate a human-annotated weight map. Row 2 presents both the weight map and the heatmap computed using Grad-CAM, showing that the two are largely consistent.
S2.5 Implementation Details
Following [28], we initialize the weights of $\mathcal{E}$, $\mathcal{D}$, and $\epsilon_{\theta}$ with the pre-trained Ip2p [3] weights. Throughout the training process, we maintain the fixed parameters of $\mathcal{E}$ and $\mathcal{D}$, focusing on training $\tau_{\theta}$ and $\epsilon_{\theta}$. The frozen emotion predictor $\mathcal{P}$ is only used during inference. We conduct experiments with NVIDIA RTX A5000 GPUs, implementing our PyTorch-based framework using the ADAM optimizer. The max time step $T$ is set to 1,000.
Figure S9: Preference of Our Method Over SOTA Methods. The first row shows the overall results, the second row only considers the trials with transitions from negative to positive emotions, and the third row focuses on the trials with transitions from positive to negative emotions. Chance is 50% in the red dotted line. Error bars are standard errors.
@@
We assess the generated outputs of all methods using Amazon Mechanical Turk (MTurk)[26] (Online). We recruit 136 participants, with each participant undergoing 180 trials, yielding a total of 24,480 trials. Figure S9 illustrates the preference of our method over other state-of-the-art methods, demonstrating the superiority of our EmoEditor in producing results that evoke the target emotion compared to other state-of-the-art methods.
From the overall results (Row 1), we see that human participants consistently prefer the generated results of our EmoEditor over all competitive baselines. Results from the “To positive" (Row 2) and “To negative" (Row 3) indicate that Neural-Style-Transfer (NST) [6] and CLIP-Styler (Csty) [13] perform comparably to our method in generating images that evoke negative emotions, but struggle with positive outcomes. This indicates that negative emotions are easily influenced by factors such as chaotic textures and overall tones, whereas positive emotions are more challenging and require an understanding of the source image along with effective adjustments. Our EmoEditor significantly outperformed Color Transfer (CT) [20] and AIF [30] in all scenarios. This confirms that emotion-evoked image generation involves more than just global color and brightness adjustments. Additionally, our EmoEditor surpasses SDEdit [16] and Ip2p [3] in generating results evoking positive emotions, while the improvement in generation performance is even more significant in generating results evoking negative emotions. This highlights challenges in current diffusion models for understanding and generating emotion-evoked images and indirectly confirms the greater challenge of generating positive emotion-evoked images compared to negative ones.
For both positive and negative emotion generation, our method is preferred over EmoEdit in human evaluations. Furthermore, the performance of our EmoEditor and the Large Model Series (LMS) is comparable. While LMS achieves slightly higher scores in positive emotion generation, the margin is narrow and largely influenced by its reliance on powerful external reasoning modules. In contrast, our method offers a compact and fully end-to-end solution that generalizes well across both positive and negative emotions, without the need for multi-stage processing or large-scale language models. Moreover, these language models are often proprietary and equipped with safety constraints, which limit their applicability in privacy-sensitive contexts.
@@
In addition, we show the confusion matrix of different methods in Figure S10. The confusion matrix provides more detailed insights into the emotion transition effects of our EmoEditor across specific emotion categories. From the upper quadrant of the confusion matrix, it can be observed that people prefer Neural-Style-Transfer (NST) [6] and CLIP-Styler (Csty) [13] more in the manipulation of emotions from positive to negative. This preference may arise because these methods excel in inducing negative emotions through distortions and irregular textures on the source image. In contrast, our EmoEditor achieves higher human preference scores in the lower quadrant of the confusion matrix, demonstrating our method’s significant superiority in generating positive-valence images. Overall, for generating positive-valence images, our EmoEditor achieves the best results in generating images that evoke awe and contentment, but it is slightly less effective in generating images that evoke excitement. This may be because excitement is a more dynamic and complex emotion, requiring a more intricate combination of visual elements.
@@
Figure S10: Confusion matrix of different methods. The confusion matrix reflects the human preference for our EmoEditor over other methods based on source and target emotion pairs. The top-left matrix presents the average overall results of our EmoEditor compared to the other SOTA methods. The remaining nine matrices show individual comparisons between our EmoEditor and each of the nine SOTA methods. Average preferences for our EmoEditor and other methods are indicated above each matrix. The vertical axis corresponds to the source emotion, while the horizontal axis represents the target emotion. Each cell in the matrix denotes the ratio at which our method is selected for the corresponding transition from the source to the target emotion.
Table S1 presents the quantitative evaluation of generated images for all competitive methods. Our EmoEditor achieves the best performance in both $S_{emo}$ and ESMI. Although Ip2p scores highest on $S_{str}$, this is because it makes minimal changes to the original images in most cases, as reflected by its low $S_{emo}$ scores and visual results. In contrast, our method effectively edits the emotion-evoking regions in the source images, achieving a balance between evoking the target emotion and preserving the structure, which is supported by the highest ESMI score.
Metric CT NeurST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
$S_{emo}$(%)↑ 27.70 65.50 38.54 36.34 39.16 19.92 31.42 46.88 38.77 69.89
$S_{str}$(%)↑ 28.96 29.47 30.11 29.55 32.65 34.82 28.00 31.60 32.03 33.22
ESMI(%)↑ 28.33 47.48 34.33 32.94 35.91 27.37 29.71 39.24 35.40 51.56
Table S1: Quantitative Evaluation of Generated Images for All Competitive Methods. Best in bold. Larger (↑) is better.
Figure S11: Visualization of Generated Images from Different Methods (from negative to positive). The target emotion is highlighted in red and the source image is framed in green.
Figure S12: Visualization of Generated Images from Different Methods (from positive to negative). The target emotion is highlighted in red and the source image is framed in green.
@@
Figure S11 illustrates the visualizations of images generated by all methods for the task of generating positive-valence images from negative emotions. Color-Transfer (CT) [20] proves ineffective for significant emotional enhancement because it primarily replicates the color features of the reference image. For instance, in example(b), it enhances brightness but fails to evoke awe as it does not alter elements causing fear. Neural-Style-Transfer (NST) [6] heavily relies on randomly selected reference images, lacking in preserving the fundamental semantic content of the source image. For example, in example(a), it generates multiple differently colored blocks but fails to evoke the emotion of excitement. CLIPstyler (Csty) [13], SDEdit [16], and CtrlNet [38] often introduce inexplicable textures, losing the semantics and structure of the source images.
Ip2p [3], due to its limited understanding of emotions, often produces results identical to the source image, showcasing restricted image generation capabilities. AIF [30] heavily relies on text descriptions for understanding emotions and can only perform global edits, similar to applying a filter. As a result, it fails to achieve effective local edits to evoke the target emotion. EmoEdit [34] has difficulty accurately identifying and localizing strongly emotion-evoking regions in source images, resulting in ineffective or imprecise edits. In example (a), Ip2p, AIF, and EmoEdit fail to understand that the fear in the source image is caused by a strange shadow. As a result, they do not alter the shadow and thus cannot counteract the fear. Even with a series of large models in LMS, it does not consistently generate ideal outputs to trigger the target emotion. In addition, it often greatly changes the scene of the source image to achieve the goal of changing the emotion. In example (c), it fails to recognize that the large flames in the background are the primary source of the original emotion, and instead generates an image where the person remains in front of the fire, merely adjusting the flame colors to be more vibrant. Such changes are insufficient to alter the original emotional impact.
In contrast, our EmoEditor, requiring only the target emotion and source image inputs, can generate highly creative images that evoke the target emotion while striving to maintain scene structure and semantic coherence. In example(a), our EmoEditor understands that the shadow in the source image is the key factor triggering fear. Instead of altering the shadow’s shape, it generates a sunset sky background, transforming the shadow into a cloud in the sky, which successfully evokes the target emotion of excitement. In example(b), it replaces a Halloween pumpkin with a castle to elicit the target emotion of awe. In example(c), it replaces the background flames with a red sunset in an attempt to evoke contentment. Notably, the generation process involves no manual intervention, showcasing the imagination and creativity of our model.
Additionally, Figure S12 shows the generation results of different methods for transitioning from positive to negative emotions. Compared to generating positive-valence images, CT and AIF perform better in generating negative-valence images. This indicates that negative emotions are more easily evoked than positive ones and can be achieved through global edits to the source image.
In addition to its creativity in generating positive-valence images, our EmoEditor also exhibits strong understanding and editing capabilities in generating negative emotions. In example(a), EmoEditor not only darkens the image’s tone to create a fearful atmosphere but also transforms the field into a desolate scene with weeds and branches, effectively evoking the emotion of fear. In example(c), it submerges the person, previously relaxing on a swim ring, beneath the pool water and alters his appearance to look frail and distressed, resembling a drowning victim, to evoke sadness. This highlights that EmoEditor can comprehend the content of the source image, align with its original emotional state, and combine it with the target emotion to perform specific edits, addressing both global and local aspects.
S3.3 CrossTest Evaluation Results
Our method employs an emotion predictor to guide emotional expression during image editing. However, using the same type of predictor for both generation and evaluation can introduce bias, as models may exploit specific predictor behaviors to inflate performance. To mitigate this, we split EmoSet into two disjoint subsets and trained two independent emotion predictors: one for generation and one for evaluation.
As shown in Table S2, NST achieves the highest ESMI score. However, as illustrated in our visualization results, its outputs often feature emotion-colored overlays that obscure the original image content. These artifacts tend to inflate $S_{\text{emo}}$ because the emotion predictor is overly sensitive to low-level color cues rather than genuine emotional semantics. This exposes a key limitation of current evaluation practices and underscores the need for more semantically grounded emotion understanding.
In contrast, results from our human study (Figure S9) show a clear preference for our method. Participants favored images where emotion-evoking regions were meaningfully altered, rather than simply covered with ambiguous color patches. This suggests that semantic coherence in emotional edits is more impactful and preferable than superficial visual modifications.
Our method ranks second, only 2% below NST, while outperforming all other baselines. Unlike NST, it preserves visual semantics while achieving strong emotional alignment, demonstrating robust generalization and a more balanced trade-off between content fidelity and emotional expressiveness.
ESMI(%) ↑ CT NST Csty CtrlNet SDEdit Ip2p AIF EmoEdit LMS Ours
CrossTest 28.39 45.35 32.61 39.38 38.50 26.54 28.67 38.50 34.09 43.95
Table S2: Quantitative Comparison of Competing Methods under CrossTest. CrossTest refers to the evaluation scheme where images are generated using an emotion predictor trained on the EmoSet subset 1 and tested by an emotion predictor trained on the EmoSet subset 2. Bold indicates the best performance, while the second-best is underlined. Larger (↑) is better.
@@
While EmoEditor excels in quantitative performance and visual results, it encounters difficulties in handling fine details of individuals, especially when their faces in the images are small. For example, in Figure S13, although our EmoEditor attempts to make the boy appear sad, the boy’s face seems somewhat twisted and distorted.
Moreover, although we use the 8-category Mikels model, real emotions are far more diverse and complex, and distinguishing positive emotions is often more challenging than negative ones. This is an area for future exploration.
Figure S13: Failure Case. The target emotion is highlighted in red and the source image is framed in green.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its NRFF award NRF-NRFF15-2023-0001 and Mengmi Zhang’s Startup Grant from Nanyang Technological University.